How would you solve the following system by elimination:
11x - 10y = -18
x - 7y = -26
Thank you.


Answer by b_souj(3)

To solve the following system by elimination:
11x - 10y = -18
x - 7y = -26
Match the coefficients of the variable in the given equations by multiplying a constant
eg.,
11x - 10y = -18
11(x - 7y = -26)
OR
7(11x - 10y = -18)
10(x - 7y = -26)
And then from the resulting equations
subtract/ add the equations that eliminates one of the variable and gives value for another variable